# UniAP: Unifying Inter- and Intra-Layer Automatic Parallelism by Mixed Integer Quadratic Programming

### v4

@@ 3.5 Complexity analysis
Let $\lvert{\mathbb{V}}\rvert$, $\lvert{\mathbb{S}}\rvert$, and $n$ denote the number of layers, parallel strategies, and GPUs, respectively. As illustrated in Algorithm 1, UniAP enumerates all factors of $n$ except 1 as the $pp\_size$ in the outer loop and all factors of $B$ except 1 in the inner loop. Since the time complexity of enumerating factors of any positive integer $a~{}(a\in{\mathbb{N}}_{+})$ is $\mathcal{O}(\sqrt{a})$, the time complexity for the two loops in the UniAP algorithm is $\mathcal{O}(\sqrt{Bd})$. Within the inner loop body, UniAP calls CostModeling to model the cost of each stage for each parallel strategy. Furthermore, the optimization time limit of the MIQP solver can be set as a constant hyperparameter when UniAP calls it. Therefore, the overall computational complexity of UniAP is $\mathcal{O}(\lvert{\mathbb{V}}\rvert\lvert{\mathbb{S}}\rvert\sqrt{Bd})$.
@@ 4 Experiment
We conduct experiments on five different kinds of environments. EnvA refers to a node (machine) with 1 Xeon 6248 CPU, 8 V100-SXM2 32GB GPUs, and 472GB memory. EnvB refers to 2 nodes interconnected with 10Gbps networks, where each node has 2 Xeon E5-2620 v4 CPUs, 4 TITAN Xp 12GB GPUs, and 125GB memory. EnvC refers to a node with 8 A100 40GB PCIe GPUs. EnvD has 4 nodes, each with the same configuration as that in EnvB. EnvE is a cloud cluster with 8 nodes and each node has 4 DCUs (a type of non-NVIDIA GPU).
We evaluate UniAP with five Transformer-based models, BERT-Huge [1], T5-Large [3], ViT-Huge [4], Swin-Huge [5], and Llama [6; 7]. We follow the common practice of training these transformer-based models. To eliminate factors that affect training throughput, we turn off techniques orthogonal to parallel strategies, such as activation checkpointing [40]. However, we integrate FP16 mixed precision training [41] for the largest model, Llama, to successfully orchestrate the process. More details on these models are provided in Appendix E.
The experimental evaluation concentrates on two primary metrics: training throughput and strategy optimization time. The former is calculated by averaging throughput from the 10th to the 60th iteration of training, while the latter is determined by measuring the time of the UOP. More details are provided in Appendix E.
Table 1: Training throughput and strategy optimization time on five Transformer-based models.
@@
EnvB BERT-Huge 6.27 $\pm$ 0.17 8.95 $\pm$ 0.06 10.77 $\pm$ 0.13 1.20 1.71
@@
EnvB BERT-Huge 2.04 $\pm$ 0.010 $>$ 40 1.51 $\pm$ 0.005 1.34 $>$ 26.32
@@ 4.1 Training throughput and strategy optimization time
We compare the training throughput and strategy optimization time of UniAP with those of the baselines on EnvA, EnvB, and EnvC. For experiments conducted on EnvA, we set the mini-batch size to be 32, 16, 128, and 128 for BERT, T5, ViT, and Swin, respectively. For experiments conducted on EnvB, we set the mini-batch size to be 16, 8, 64, and 32 for these four models, respectively. For the Llama model [6; 7] run on EnvC, we set the mini-batch size to be 8.
@@
We choose Galvatron [37] and Alpa [25] as baselines because they have achieved SOTA performance. Specifically, Galvatron has surpassed other methods, including PyTorch DDP [16], Megatron-LM [18], FSDP [42; 19], GPipe [8], and DeepSpeed 3D [43] in terms of training throughput, as reported in the original paper [37]. Furthermore, Alpa utilizes the Just-In-Time (JIT) compilation feature in JAX and outperforms Megatron-LM and DeepSpeed.
@@
Table 1 shows the training throughput and strategy optimization time on EnvA, EnvB, and EnvC. On EnvA, UniAP and Galvatron get the same optimal strategy for BERT-Huge, T5-Large, and ViT-Huge, outperforming Alpa in terms of training throughput and strategy optimization time. In addition, UniAP finds a solution for Swin-Huge, while Galvatron encounters CUDA OOM problems. In particular, UniAP achieves a maximum optimization speedup that is 17$\times$ faster than Galvatron and hundreds of times faster than Alpa on BERT-Huge. This is mainly due to the ability of the MIQP solver to search for an optimal strategy on multiple threads, while the dynamic programming based methods like Galvatron and Alpa run on a single thread due to their strong data dependency.
@@
On EnvB, UniAP consistently demonstrates competitive or larger training throughput compared to Galvatron and Alpa. We attribute the performance improvement to the larger strategy space in UniAP. A detailed study of the importance of strategy space is provided in Section 4.4. Furthermore, the strategy optimization time of UniAP is also significantly shorter than the two baseline methods.
@@
On EnvC, UniAP shows an optimization speedup of 11.83$\times$ and a training speedup of 3.80$\times$ compared to Galvatron on Llama. We examine the parallel strategy they adopted. The results show that UniAP employs an 8-stage PP with a micro-batch size of 1, whereas Galvatron employs a 4-stage PP with a micro-batch size of 3 (with the final micro-batch containing 2 samples). Within each PP stage, Galvatron utilizes a 2-way TP. For EnvC, which is equipped with 8 A100 40GB PCIe GPUs, minimizing communication volume is critical. Since PP has significantly less inter-device communication volume than TP, the parallel strategy discovered by UniAP is more reasonable than Galvatron in this case.
@@
To facilitate further discussions, we provide another case study of the optimal parallel strategy for BERT-Huge in Appendix F.
4.2 Estimation accuracy
Some variables in UniAP and other AP methods are estimated values rather than actual running values. The TPI (inverse of training throughput) is the most important one. Accurate estimation for TPI or training throughput is crucial for evaluating candidate parallel strategies and ensuring the optimality of the solution. We hereby introduce a metric called relative estimation error (REE) $e$ to quantify the accuracy of the estimated training throughput $\hat{T}$ to the actual training throughput $T$:
$e(T,\hat{T})=\frac{|T-\hat{T}|}{T}\times 100\%.$ (9)
We evaluate the optimal parallel strategies obtained from EnvA and EnvB. Results show that UniAP achieves an average REE of 3.59%, which is lower than 11.17% for Galvatron.
4.3 Scalability
@@
We study the scalability of UniAP using BERT, T5, ViT, and Swin on EnvD. We set the mini-batch sizes for each data point as 8, 4, 32, and 16 times the number of nodes (denoted as ‘#nodes’). The experimental results are shown in Figure 4. In Figure 4(a), the training throughput of the optimal strategy demonstrate near-linearity as the number of nodes and mini-batch size increase. In Figure 4(b), the strategy optimization time matches the complexity analysis in Section 3.5.
Additionally, we use EnvE to scale UniAP to larger clusters and larger models including Llama-13B. UniAP also outperforms other baselines in this experiment. Details are available in Appendix G.
4.4 Ablation study
Table 2: Ablation study on the importance of unifying strategy space.
Model Training throughput (samples/s)
UniAP (Inter-only) UniAP (Intra-only) UniAP
BERT-Huge SOL×$\times$1) 2.48 $\pm$ 0.02 10.77 $\pm$ 0.13
T5-Large SOL×$\times$1 2.92 $\pm$ 0.01 9.01 $\pm$ 0.06
ViT-Huge 45.58 $\pm$ 0.54 CUDA×$\times$2 45.58 $\pm$ 0.54
Swin-Huge 19.08 $\pm$ 0.10 4.66 $\pm$ 0.02 19.08 $\pm$ 0.10
  • 1)
    SOL×$\times$: No solution after strategy optimization.
  • 2)
    CUDA×$\times$: CUDA OOM exceptions during model training.
We investigate the importance of the strategy space for the optimality of parallel strategies with an ablation study. Specifically, we constrain the strategy space to inter-layer-only and intra-layer-only strategies and evaluate the training throughput of the resulting optimal strategy on EnvB. We set the mini-batch sizes to be 16, 12, 64, and 32, respectively.
Results are shown in Table 2. We can find that constraining the strategy space compromises the optimality of parallel strategies or gets strategies that encounter OOM across different models. Hence, unifying inter- and intra-layer AP for joint optimization is essential and necessary.
@@ 5 Conclusion
Appendix A Visualization for parallel methods
In order to provide an intuitive analysis, we present visualizations in Figure 5 on different parallel methods for optimizing parallel strategies for a three-layer model. The numbered boxes represent the corresponding layers of the model. The layers with gray backgrounds represent the final parallel strategy for each method. The layers without dotted lines indicate that DP was applied to them, while TP was applied to the others. The direction of the dotted lines indicates the dimension of the tensor that TP is applied to. The solid arrows between layers represent connections within a single pipeline stage, while the dashed arrows represent connections between consecutive pipeline stages. For example, the parallel strategy optimized by MP (Figure 5(b)) represents that PP and TP are utilized to train the model. Specifically, layer 0 of the model is assigned to the first pipeline stage, while layer 1 and 2 are assigned to the second pipeline stage. Additionally, TP is activated across all model layers.
As shown in Figure 5(b), MP methods are manually designed and rely on expert knowledge, resulting in limited flexibility. Inter-layer-only and intra-layer-only AP methods (Figure 5(c) and 5(d)) optimize (search) from a set of candidate inter-layer-only and intra-layer-only parallel strategies, respectively. To the best of our knowledge, existing inter- and intra-layer parallelism AP methods optimize parallel strategies hierarchically. Thus, we refer to them as Hierarchical AP in Figure 5(e). They initially adopt algorithms, such as greedy or dynamic programming, to propose candidate inter-layer parallel strategies. Subsequently, they optimize the intra-layer parallel strategy for every fixed inter-layer parallel strategy. Compared to them, UniAP jointly optimize all possible parallel strategies from intra-layer-only parallelism to inter- and intra-layer parallelism. Consequently, UniAP has the largest strategy space for exploration.
Appendix B Proof of the linear form for the contiguous set
To facilitate our discussion, we adopt the linear form of the order-preserving constraint as presented in the main paper. We denote ${\boldsymbol{\mathit{P}}}_{ui}$ as a 0-1 variable indicating whether layer $u$ is to be placed on the $i$-th computation stage, $pp\_size$ as the number of computation stages in the pipeline. Besides, $\mathcal{G}({\mathbb{V}},{\mathbb{E}})$ represents the computation graph for the model. Then, we formalize the theorem as follows:
Theorem B.1.
A subgraph ${\mathbb{V}}_{i}=\{\forall u\in{\mathbb{V}}:{\boldsymbol{\mathit{P}}}_{ui}=1\}$ is contiguous if and only if there exists ${\boldsymbol{\mathit{Z}}}_{vi}$ such that Equation (6a), (6b), and (6c) are satisfied.
@@ Appendix B Proof of the linear form for the contiguous set
"If": Assume that there exists nodes $u,w\in{\mathbb{V}}_{i}$ and $v\notin{\mathbb{V}}_{i}$ such that $v$ and $w$ are reachable from $u$ and $v$, respectively. Hence, ${\boldsymbol{\mathit{P}}}_{ui}=1$, ${\boldsymbol{\mathit{P}}}_{wi}=1$, and ${\boldsymbol{\mathit{P}}}_{vi}=0$. Without losing generality, we assume $\langle u,v\rangle\in{\mathbb{E}}$. Thus, according to Equation (6c), we have ${\boldsymbol{\mathit{Z}}}_{vi}\leqslant{\boldsymbol{\mathit{P}}}_{vi}-{\boldsymbol{\mathit{P}}}_{ui}+1=0$. By applying Equation (6b) repeatedly following the path from $v$ to $w$, we have ${\boldsymbol{\mathit{Z}}}_{wi}\leqslant{\boldsymbol{\mathit{Z}}}_{vi}$. Thus, ${\boldsymbol{\mathit{Z}}}_{wi}\leqslant 0$. However, we also have ${\boldsymbol{\mathit{Z}}}_{wi}\geqslant{\boldsymbol{\mathit{P}}}_{wi}=1$ according to Equation (6a). A contradiction.
"Only if": First, we define ${\boldsymbol{\mathit{Z}}}_{vi}=1$ if a node $w\in{\mathbb{V}}_{i}$ is reachable from $v~{}(v\in{\mathbb{V}})$. Otherwise, ${\boldsymbol{\mathit{Z}}}_{vi}=0$. Thus, Equation (6a) and (6b) are satisfied according to this kind of definition. For Equation (6c), if ${\boldsymbol{\mathit{P}}}_{vi}=1$, the constraint will hold true regardless of whether ${\boldsymbol{\mathit{P}}}_{ui}$ is $1$ or 00. If ${\boldsymbol{\mathit{P}}}_{vi}=0$ and ${\boldsymbol{\mathit{P}}}_{ui}=0$, ${\boldsymbol{\mathit{Z}}}_{vi}\leqslant{\boldsymbol{\mathit{P}}}_{vi}-{\boldsymbol{\mathit{P}}}_{ui}+1=1$ will also hold true because ${\boldsymbol{\mathit{Z}}}_{vi}$ could be either 00 or $1$. Finally, if ${\boldsymbol{\mathit{P}}}_{vi}=0$ and ${\boldsymbol{\mathit{P}}}_{ui}=1$, ${\boldsymbol{\mathit{Z}}}_{vi}=0$ will hold true because ${\mathbb{V}}_{i}$ is a contiguous set and we cannot find any $w\in{\mathbb{V}}_{i}$, such that $w$ is reachable from $v$. ∎
(a) Model
(b) MP
(c) Inter-layer-only AP
(d) Intra-layer-only AP
(e) Hierarchical AP
(f) UniAP
Figure 5: Parallel methods for optimizing parallel strategies for a three-layer model in (a).
Appendix C QIP formulation for intra-layer-only parallelism

### v5

@@ 3.5 Complexity analysis
Let $\lvert\mathbb{V}\rvert$, $\lvert\mathbb{S}\rvert$, and $n$ denote the number of layers, parallel strategies, and GPUs, respectively. As illustrated in Algorithm 1, UniAP enumerates all factors of $n$ except 1 as the $deg$ in the outer loop and all factors of $B$ except 1 in the inner loop. The time complexity for the two loops in the UniAP algorithm is $\mathcal{O}(\sqrt{Bn})$. Within the inner loop body, UniAP calls CostModeling to model the cost of each stage for each parallel strategy. Furthermore, the optimization time limit of the MIQP solver can be set as a constant hyperparameter when UniAP calls it. Therefore, the overall computational complexity of UniAP is $\mathcal{O}(\lvert\mathbb{V}\rvert\lvert\mathbb{S}\rvert\sqrt{Bn})$.
@@ 4 Experiment
4.1 Experiment setup
We conduct experiments on four different kinds of environments to validate the effectiveness and universality of our method. EnvA refers to a node (machine) with 1 Xeon 6248 CPU, 8 V100-SXM2 32GB GPUs, and 472GB memory. EnvB refers to 4 nodes interconnected with 10Gbps networks and each node has 2 Xeon E5-2620 v4 CPUs, 4 TITAN Xp 12GB GPUs, and 125GB memory. EnvC refers to a node with 8 A100 40GB PCIe GPUs. EnvD is a cloud cluster with 16 nodes and each node has 4 DCUs (a type of non-NVIDIA GPU). In the following text, we will specify the number of GPUs used after the environment name (e.g., EnvB-8) only when partial nodes of the environment are used.
We evaluate UniAP with five Transformer-based models, BERT-Huge [5], T5-Large [31], ViT-Huge [6], Swin-Huge [24], and Llama [40, 41]. We follow the common practice of training these transformer-based models. To eliminate factors that affect training throughput, we turn off techniques orthogonal to parallel strategies, such as activation checkpointing [3]. However, we integrate FP16 mixed precision training [26] for the largest model, Llama, in some cases to successfully orchestrate the process. More details on these models are provided in Appendix D.
The experimental evaluation concentrates on two primary metrics: training throughput and strategy optimization time. The former is calculated by averaging throughput from the 10th to the 60th iteration of training, while the latter is determined by measuring the time of the UOP. More details are provided in Appendix D. Besides, we provide the accuracy of our cost model in Appendix E.
@@ 4.1 Experiment setup
EnvB-8 BERT-Huge 6.27 $\pm$ 0.17 8.95 $\pm$ 0.06 10.77 $\pm$ 0.13 1.20 1.71
@@
EnvB-8 BERT-Huge 2.04 $\pm$ 0.010 $>$ 40 1.51 $\pm$ 0.005 1.34 $>$ 26.32
@@
Table 1: Training throughput and strategy optimization time on EnvA, EnvB-8, and EnvC.
Megatron DeepSpeed UniAP Megatron DeepSpeed UniAP
  • 1)
    SOL×$\times$: No solution after strategy optimization.
Table 2: Training throughput and strategy optimization time on EnvD-32.
1 st.1) 2 nd.2) Slowest3) Median4)
Llama-7B 8 2.01 1.92 0.22 0.82 41 64
Llama-13B 4 0.82 0.58 0.27 0.42 42 48
  • 1)
    1 st.: The training throughput achieved by the fastest parallel strategy.
  • 2)
    2 nd.: The training throughput achieved by the second fastest parallel strategy.
  • 3)
  • 4)
  • 5)
  • 6)
    #candidate: The total number of candidate parallel strategies.
@@ 4.2 Training throughput and strategy optimization time
We compare the training throughput and strategy optimization time of UniAP with those of the baselines on EnvA, EnvB-8, EnvC, and EnvD-32. For experiments on EnvA, we set the mini-batch size to be 32, 16, 128, and 128 for BERT, T5, ViT, and Swin, respectively. For experiments on EnvB-8, we set the mini-batch size to be 16, 8, 64, and 32 for these four models, respectively. For the Llama model [40, 41] run on EnvC, we set the mini-batch size to be 8. For experiments on EnvA, EnvB-8, and EnvC, we choose Galvatron [25] and Alpa [47] as baselines because they have achieved SOTA performance. Specifically, Galvatron has surpassed other methods, including PyTorch DDP [22], Megatron-LM [30], FSDP [32, 8], GPipe [15], and DeepSpeed 3D [27] in terms of training throughput, as reported in the original paper [25]. Furthermore, Alpa utilizes the Just-In-Time (JIT) compilation feature in JAX and outperforms Megatron-LM and DeepSpeed.
For benchmarks on EnvD-32, pytorch libraries and frameworks need specialized adaptations due to the heterogeneity of DCUs against NVIDIA GPUs. Thus, this part of experiments is based on adapted pytorch libraries and Megatron-DeepSpeed[27] framework provided by the platform developers, and we select MP methods including Megatron [30] and DeepSpeed (ZeRO-3) [34] as our baseline methods. We set the mini-batch size as 8 for Llama-7B and 4 for Llama-13B. The remaining configurations align with those for Llama-7B in Section 4.1, such as activating FP16 mixed precision training.
@@
Table 1 shows the training throughput and strategy optimization time on EnvA, EnvB-8, and EnvC. On EnvA, UniAP and Galvatron get the same optimal strategy for BERT-Huge, T5-Large, and ViT-Huge, outperforming Alpa in terms of training throughput and strategy optimization time. In addition, UniAP finds a solution for Swin-Huge, while Galvatron encounters CUDA OOM issues. In particular, UniAP achieves a maximum optimization speedup that is 17$\times$ faster than Galvatron and hundreds of times faster than Alpa on BERT-Huge. This is mainly due to the ability of the MIQP solver to search for an optimal strategy on multiple threads, while the dynamic programming based methods like Galvatron and Alpa run on a single thread due to their strong data dependency.
@@
On EnvB-8, UniAP consistently demonstrates competitive or larger training throughput compared to Galvatron and Alpa. We attribute the performance improvement to the larger strategy space in UniAP. A detailed study is provided in Section 4.4. Furthermore, the strategy optimization time of UniAP is also significantly shorter than the two baseline methods.
@@
On EnvC, UniAP shows an optimization speedup of 11.83$\times$ and a training speedup of 3.80$\times$ compared to Galvatron on Llama. We examine the parallel strategy they adopted. UniAP employs an 8-stage PP with a micro-batch size of 1, whereas Galvatron employs a 4-stage PP with a micro-batch size of 3 (with the final micro-batch containing 2 samples). Within each PP stage, Galvatron utilizes a 2-way TP. For EnvC, which is equipped with 8 A100 40GB PCIe GPUs, minimizing communication volume is critical. Since PP has significantly less inter-device communication volume than TP, the parallel strategy discovered by UniAP is more reasonable than Galvatron in this case.
Table 2 shows the results on EnvD-32. On this environment, UniAP consistently identifies the fastest parallel strategies while expending considerably less time on strategy optimization compared to Megatron (about 157$\times$ faster for Llama-7B and 77$\times$ faster for Llama-13B). Please note that the strategy optimization time Megatron needs for Llama-7B surpasses that for Llama-13B. We attribute this phenomenon to the variations in the mini-batch size. Specifically, the escalation from a mini-batch size of 4 (employed in Llama-13B) to 8 (employed in Llama-7B) leads to a rise in the number of candidate parallel strategies, as well as the parallel strategies that will successfully train the model. As a result, the strategy optimization time will become increasingly long when the mini-batch size exceeds 8 in the pretraining scenario [1, 40, 41]. In this situation, UniAP will identify the optimal parallel strategy much faster.
@@
To facilitate further discussions, we provide a case study of the optimal parallel strategy in Appendix F.
@@
Number of Training throughput Strategy optimization
DCUs (samples/s) time(min.)
16 0.1726 $\pm$ 0.0005 0.75 $\pm$ 0.0100
32 0.3393 $\pm$ 0.0004 1.64 $\pm$ 0.0321
64 0.6656 $\pm$ 0.0023 1.55 $\pm$ 0.0058
Table 4: Scalability on training throughput among DCU clusters.
4.3 Scalability
We study the scalability of UniAP using BERT, T5, ViT, and Swin on EnvB. We set the mini-batch sizes for each data point as 8, 4, 32, and 16 times the number of nodes (denoted as ‘#nodes’). The experimental results are shown in Figure 5. In Figure 5(a), the training throughput of the optimal strategy demonstrate near-linearity as the number of nodes and mini-batch size increase. In Figure 5(b), the strategy optimization time matches the complexity analysis in Section 3.5.
Additionally, we use DCU nodes in EnvD to scale UniAP to larger clusters on Llama-7B, with precision of FP32. We conducted experiments on clusters with 16, 32, and 64 DCUs, where we set the mini-batch size to 1/8 of the number of DCUs (i.e., for a cluster with 16 DCUs, we set the mini-batch size to 2). The experimental results are shown in Table 4, further demonstrating the near-linear scalability of UniAP.
4.4 Further study on the significance of UniAP
Finally, we investigate the significance of UniAP by answering two questions. The first question is why we need AP rather than MP. We examine the statistics on the candidate parallel strategies for Megatron, as shown in Table 3. With MP, most users are unfamiliar with the complex hyper-parameter detail of the candidate parallel strategies and typically choose a parallel strategy at random, which will result in a 64.1% (41/64) chance of failing to train Llama-7B and an 87.5% (42/48) chance of failing to train Llama-13B even when the hardware resources are sufficient. Even if they successfully eliminate all infeasible parallel strategies, they still only have a 50% chance of identifying a parallel strategy with training throughput better than 0.82 samples/s for Llama-7B and 0.42 samples/s for Llama-13B. When the users can’t identify the fastest parallel strategy out of hundreds of candidates, they may sacrifice at least 4.4% of the training throughput for Llama-7B and 29.2% for Llama-13B. In such circumstances, UniAP is able to identify the fastest strategy, as shown in Table 2.
Figure 6: Ablation study. SOL×$\times$ indicates no feasible solution during strategy optimization, and CUDA×$\times$ indicates CUDA OOM exceptions during training.
The second question is why UniAP achieves better performance than other AP methods. We investigate the importance of the strategy space for the optimality of parallel strategies with an ablation study. Specifically, we constrain the strategy space to inter-layer-only and intra-layer-only strategies and evaluate the training throughput of the resulting optimal strategy on EnvB. We set the mini-batch sizes to be 16, 12, 64, and 32, respectively. Results are shown in Figure 6. We can find that constraining the strategy space compromises the optimality of parallel strategies or gets strategies that encounter OOM across different models. Hence, unifying inter- and intra-layer AP for joint optimization makes UniAP outperform other AP methods.
@@ 5 Conclusion
Appendix A Proof of the linear form for the contiguous set
To facilitate our discussion, we adopt the linear form of the order-preserving constraint as presented in the main paper. We denote $\textbf{{P}}_{ui}$ as a 0-1 variable indicating whether layer $u$ is to be placed on the $i$-th computation stage, $pp\_size$ as the number of computation stages in the pipeline. Besides, $\mathcal{G}(\mathbb{V},\mathbb{E})$ represents the computation graph for the model. Then, we formalize the theorem as follows:
Theorem A.1.
A subgraph $\mathbb{V}_{i}=\{\forall u\in\mathbb{V}:\textbf{{P}}_{ui}=1\}$ is contiguous if and only if there exists $\textbf{{Z}}_{vi}$ such that Equation (6a), (6b), and (6c) are satisfied.
@@ Appendix A Proof of the linear form for the contiguous set
"If": Assume that there exists nodes $u,w\in\mathbb{V}_{i}$ and $v\notin\mathbb{V}_{i}$ such that $v$ and $w$ are reachable from $u$ and $v$, respectively. Hence, $\textbf{{P}}_{ui}=1$, $\textbf{{P}}_{wi}=1$, and $\textbf{{P}}_{vi}=0$. Without losing generality, we assume $\langle u,v\rangle\in\mathbb{E}$. Thus, according to Equation (6c), we have $\textbf{{Z}}_{vi}\leqslant\textbf{{P}}_{vi}-\textbf{{P}}_{ui}+1=0$. By applying Equation (6b) repeatedly following the path from $v$ to $w$, we have $\textbf{{Z}}_{wi}\leqslant\textbf{{Z}}_{vi}$. Thus, $\textbf{{Z}}_{wi}\leqslant 0$. However, we also have $\textbf{{Z}}_{wi}\geqslant\textbf{{P}}_{wi}=1$ according to Equation (6a). A contradiction.
"Only if": First, we define $\textbf{{Z}}_{vi}=1$ if a node $w\in\mathbb{V}_{i}$ is reachable from $v~{}(v\in\mathbb{V})$. Otherwise, $\textbf{{Z}}_{vi}=0$. Thus, Equation (6a) and (6b) are satisfied according to this kind of definition. For Equation (6c), if $\textbf{{P}}_{vi}=1$, the constraint will hold true regardless of whether $\textbf{{P}}_{ui}$ is $1$ or 00. If $\textbf{{P}}_{vi}=0$ and $\textbf{{P}}_{ui}=0$, $\textbf{{Z}}_{vi}\leqslant\textbf{{P}}_{vi}-\textbf{{P}}_{ui}+1=1$ will also hold true because $\textbf{{Z}}_{vi}$ could be either 00 or $1$. Finally, if $\textbf{{P}}_{vi}=0$ and $\textbf{{P}}_{ui}=1$, $\textbf{{Z}}_{vi}=0$ will hold true because $\mathbb{V}_{i}$ is a contiguous set and we cannot find any $w\in\mathbb{V}_{i}$, such that $w$ is reachable from $v$. ∎
Appendix B QIP formulation for intra-layer-only parallelism